Following are the income statement and some additional information for Carolina Consulting Company.
  

Carolina Consulting Company
Income Statement
For the Year Ended December 31, 2018
Net sales $ 17,000
Cost of goods sold (2,900 )
Gross margin 14,100
Operating expenses $ 3,400
Depreciation expense 1,600 (5,000 )
Income before taxes 9,100
Income taxes (3,000 )
Net income $ 6,100

  
All sales were on credit and accounts receivable decreased by $1,040 in 2018 compared to 2017. Merchandise purchases were on credit with a decrease in accounts payable of $840 during the year. Ending inventory was $640 larger than beginning inventory. Income taxes payable increased $440 during the year. All operating expenses were paid for in cash.
  
Required:
Prepare the cash flows from operating activities section of the statement of cash flows using the direct method. (Amounts to be deducted should be indicated with a minus sign.)

Cash Flows from Operating Activities: Net cash flows from operating activities
Answer #1
Cash flows from operating activities:
Cash received from customers 18040 =17000+1040
Cash paid to suppliers -4380 =2900+840+640
Cash paid for operating expenses -3400
Cash paid for taxes -2560 =3000-440
Net cash flows from operating activities 7700